# The Landscape of College-level Data Visualization Courses, and the Benefits of Incorporating Statistical Thinking

### v1

@@ Abstract
Data visualization is a core part of statistical practice and is ubiquitous in many fields. Although there are numerous books on data visualization, instructors in statistics and data science may be unsure how to teach data visualization, because it is such a broad discipline. To give guidance on teaching data visualization from a statistical perspective, we make two contributions. First, we conduct a survey of data visualization courses at top colleges and universities in the United States, in order to understand the landscape of data visualization courses. We find that most courses are not taught by statistics and data science departments and do not focus on statistical topics, especially those related to inference. Instead, most courses focus on visual storytelling, aesthetic design, dashboard design, and other topics specialized for other disciplines. Second, we outline three teaching principles for incorporating statistical inference in data visualization courses, and provide several examples that demonstrate how instructors can follow these principles. The dataset from our survey allows others to explore the diversity of data visualization courses, and our teaching principles give guidance to instructors and departments who want to encourage statistical thinking via data visualization. In this way, statistics-related departments can provide a valuable perspective on data visualization that is unique to current course offerings.
@@ 1 Introduction: How is Data Visualization Taught, and What Defining Role Can Statistics Play?
Nearly every field involves data visualization, where information from data are visualized in order to gain insight on some phenomena. Data visualizations can be as simple as a bar plot of categorical counts to as complex as an interactive dashboard involving maps, time series, networks, and other high-dimensional graphs. Given its ubiquity, data visualization is taught across many fields, each with their own unique goals. As a result, there are many books dedicated to data visualization within specific fields, including business (Soukup and Davidson,, 2002; Knaflic,, 2015; Zheng,, 2017), journalism (Cairo,, 2012), and the humanities (Engebretsen and Kennedy,, 2020). There are also many books about data visualization in general (Tufte,, 2001; Chen et al.,, 2007; Wilke,, 2019; Cairo,, 2016, 2019, 2023; Yau,, 2024) and data visualization software (Sarkar,, 2008; Jones,, 2014; Murray,, 2017; Sievert,, 2020). Data visualization is such a pervasive and diverse area that, perhaps paradoxically, instructors in statistics, data science, and related fields may be unsure how to teach data visualization.
Indeed, the importance of data visualization in the practice and education of statistics has been emphasized for decades. For example, in the 2016 Guidelines for Assessment and Instruction in Statistics Education (GAISE) College Report endorsed by the American Statistical Association, one of the 10 goals is that students “should be able to produce graphical displays and numerical summaries and interpret what graphs do and do not reveal” (Carver et al.,, 2016). Although data visualization is likely used in most (if not all) introductory statistics courses, instructors and departments may wonder if there is value in teaching data visualization specifically, perhaps even in a standalone course. For example, our statistics and data science department at Carnegie Mellon University offers data visualization courses at the undergraduate and graduate level. And even if departments do not plan to dedicate a semester-long class to data visualization, they may wonder how they can incorporate instruction of data visualization into existing courses.
@@ 2. Question 2: What defining role can statistics play in teaching data visualization?
To answer the first question, we collected data on 135 highly ranked colleges and universities, and determined which courses on data visualization, if any, were offered at each school in the 2022-2023 academic year. Among these 135 schools, we identified 270 data visualization courses. We recorded the name of each course, whether it was an undergraduate- and/or graduate-level course, the department(s) that offered it, and the topics and software taught based on its course description. There were three notable conclusions from our search. First, many schools offered data visualization courses (94 out of 135, or 69.6%), and often several departments within a school offered their own data visualization courses. The most frequent departments were related to computer science, information systems, management, and analytics, but we also found courses in less stereotypically data-centric fields, like geography, journalism, and the visual arts. Second, we found that most data visualization courses were not taught in statistics and data science departments (220 out of 270, or 81.5%), which may come as a surprise to some readers. We found that while data visualization is frequently taught as one component of statistics courses (e.g., exploratory data analysis in introductory statistics courses), it is also rarely the focus of the course. Finally, we found that most data visualization courses did not focus on statistical topics, and instead tended to specialize in visual storytelling, aesthetic design, or software use and development. In particular, the vast majority of courses did not focus on key aspects of statistical thinking: testing, uncertainty quantification, and modeling. For example, although many courses taught students how to visually examine trends and patterns in data, they rarely taught students how to visualize uncertainty in these trends or determine if these were “statistically significant” findings.
@@
To demonstrate how statistics departments can do this, we discuss how data visualization is taught in our statistics and data science department at Carnegie Mellon University. When our department first started teaching data visualization, it focused on visual design choices and how to use statistical software, which are common topics covered by many other college-level courses, as revealed by our survey. But over time, our course has evolved to put more emphasis on statistical issues that arise when creating data visualizations. For example, how should students pinpoint whether a graphical finding is “statistically significant,” and how can they quantify their uncertainty about trends and patterns they visualize? Furthermore, if multiple comparisons can be made within a single graph, how should the visual account for these multiple comparisons in order for inference to be correct? Addressing these questions forced us to develop a taxonomy of graphs and analyses, such that students can better understand how inferential tools can be properly used to complement their graphs, and vice versa. Our course isn’t the only one that emphasizes statistical aspects of data visualization; in our survey, we found several other courses that emphasize similar topics. Nonetheless, these courses are by far the minority among data visualization courses.
@@ 2.1 Description of Survey
For each university and college, we searched their online course catalog for credit-granting courses in Fall 2022 or Spring 2023 that contained one or more of the following words or combinations of words: (1) data visualization, (2) visualization / visualize / visualizing, (3) graph(s) / graphics / graphing, (4) plot(s) / plotting, and/or (5) chart(s) / charting. After identifying all courses that contained these words, we carefully reviewed the course titles, descriptions, and syllabi or websites (where available), and declared a course as a data visualization course only if its primary focus was indeed data visualization. For instance, there are many introductory statistics and data science courses that cover data visualization in some capacity, but often data visualization is not the main topic of these courses, and thus we typically did not consider them data visualization courses in our survey. For simplicity, we excluded courses in professional programs (medical school, nursing school, law school, and MBA programs), such that we focus on general undergraduate and graduate programs. Additionally, we only considered a course a data visualization course if it had a data component, and not just a visualization component. For example, there are graphics courses that focus on creating video game graphics, which is a different context than visual representations of data that we are interested in.
@@
For each course that fit our search criteria, we gathered several pieces of information relevant to understanding the different contexts in which data visualization is taught. This includes the academic level (undergraduate, graduate, or both) and department in which the course was taught, as well as what software was used if listed in the course description or available syllabus/website. To additionally gauge the level of statistical content included in data visualization courses, we tracked if any of the following topics were mentioned in the course’s description, syllabus, or website (if available): hypothesis testing, confidence intervals, statistical modeling (e.g., linear regression), spatial data, clustering, time series, high dimensional analyses (e.g., principal component analysis), text data, networks, and interactive graphics (e.g., Shiny, D3.js).
@@
All replication materials for our survey results, including the dataset of data visualization courses we identified in our survey, is available at github.com/ryurko/teaching-data-viz. In what follows, we present several results about this dataset, thereby describing the current landscape of data visualization courses at U.S. colleges and universities. We hope that this dataset can also serve as a useful resource for instructors who want to glean further insights about data visualization courses, or use it as an example dataset in their own courses.
@@ 2.2 Survey Results
In total, we identified 270 data visualization courses across this pool of universities/colleges. Figure 1 displays the number of data visualization courses across schools; most schools have fewer than five courses, with a right-skew denoting places that offer several courses. Georgetown University (15 courses), New York University (14), and Northeastern University (13) stand out with the highest number.
@@
Figure 1: Histogram of the number of data visualization classes taught within a school.
@@
One of our main interests is understanding how many data visualization courses are taught by statistics and data science departments. Among the 270 courses, there are 163 unique department names, making it difficult to categorize courses by department. Indeed, departments ranged from quantitative fields (e.g., computer science, economics, engineering, mathematics, statistics) to social sciences (e.g., education, psychology, sociology) to professional domains (e.g., analytics, business, government, journalism) to humanities and the arts (e.g., anthropology, design, history, visual arts, writing), demonstrating the ubiquity of data visualization. Figure 2 shows a word cloud of the most common words across department names among these 270 courses; broadly speaking, the most common departments are related to computer science, information systems, management, and analytics. We also see many departments related to business, design, engineering, public policy, and statistics. Finally, the diversity of less common departments is notable, including the arts, biology, geography, health, journalism, and mathematics.
@@
To understand the involvement of statistics and/or data science departments specifically, we identified department names that included either “stat” or “data” in their name. [^2] The only departments that mentioned “data” but not “data science” mentioned “data analytics;” the only exception was a department called “Data for Political Research.” In order to take an inclusive view of data science, we also considered these as data science departments. We found 50 of the 270 courses (18.5%) were taught by statistics and/or data science departments. This proportion was also fairly consistent across the level of courses: Figure 3 displays the number of data visualization courses at the undergraduate or graduate level, or both (e.g. cross-listed courses), sorted by whether the course’s department was in statistics or data science. A roughly equal number of data visualization courses are offered at the undergraduate or graduate level, where 21.7% of undergraduate courses, 18.8% of graduate courses, and 12.7% of cross-listed courses are taught by statistics or data science departments. Taken together, Figures 2 and 3 demonstrate that many fields teach data visualization, most of which are outside of statistics and data science.
@@
Figure 2: Word cloud displaying the top words in the department names for the 270 identified data visualization courses.
@@
Meanwhile, Figure 4(b) shows the number of topics that are taught across courses, and we see that the vast majority of courses (226, or 88.3%) teach up to two of these 10 topics. For example, we found that many courses that taught interactive graphics focused exclusively on how to create visualization dashboards, and many courses that taught spatial data focused exclusively on how to make data-based maps with geospatial software. The only course that covers nine topics is our own course at Carnegie Mellon University (which we discuss in the next section), and the only course that covers eight topics is from Colby College. [^3] The only topic our course doesn’t cover is networks. Notably, the instructor who developed the Colby College course was previously a PhD student in our department and taught our data visualization course. However, we want to emphasize that this does not suggest that courses are inadequate by not covering these topics; many courses are certainly teaching other useful content, e.g. visual storytelling and design. Rather, departments understandably focus on topics most related to their own fields, instead of teaching statistical topics, especially those related to inference.
@@
We acknowledge that there are limitations with this survey. First, there are likely other data visualization courses not captured by our survey (e.g., at other schools not considered here, or courses that did not mention keywords we searched for). Similarly, some courses may cover statistical topics but did not mention them on publicly available sources we could access. Finally, this survey does not capture all possible topics taught in data visualization courses, because we focused on statistical topics. Nonetheless, this survey demonstrates that data visualization is taught across many colleges and universities, and by a broad range of disciplines that are largely outside of statistics or data science. Furthermore, most data visualization courses do not emphasize key aspects of statistical thinking (testing, uncertainty quantification, and modeling), and instead focus on topics unique to particular disciplines (e.g. storytelling, visual design, dashboard design, or specialized software). Thus, to complement current offerings in data visualization, in the following section we demonstrate how instructors can incorporate statistical thinking when teaching data visualization.
@@ 3.1 Description of Our Course, and a Taxonomy of Graphs
Our undergraduate data visualization course in the Department of Statistics and Data Science at Carnegie Mellon University (CMU), titled Statistical Graphics and Visualization, has existed for almost 20 years and has evolved over time. Currently, the course is offered every fall and spring semester, and serves approximately 100 students each semester. The prerequisite is two semesters of introductory statistics (not calculus-based), such that students have a basic understanding of hypothesis testing, statistical models (including linear regression), and how to apply them to real datasets. We use R to implement graphs and analyses, but students are not required to have prior programming experience. The course also counts towards our college’s general education requirement, specifically the “design” requirement, which consists of courses focused on designing products to solve problems. Our department is within CMU’s Dietrich College of Humanities and Social Sciences, and we regularly teach majors in economics, English, history, information systems, psychology, social and decision sciences, and other fields. Thus, our course is meant for a broad group of students from many disciplines who are at a sophomore-level or above in terms of their statistical background.
@@
Histograms, boxplots,
smoothed densities, empirical CDFs
@@
We’ve found that there are several pedagogical benefits to organizing the first half of the course with the taxonomy in Table 1. First, students are typically quite good at determining whether a variable is categorical or quantitative, but they may find it challenging to determine which graphs or analyses are most appropriate for a given dataset. Thus, organizing by variable type gives students a straightforward scaffolding for navigating graphs and analyses. Second, it prepares students to make and interpret more complex graphs in the second half of the course (e.g., high-dimensional visualizations, maps, interactive graphs), which frequently use the graphs in Table 1 as building blocks. Third, the taxonomy encourages a spiral learning approach in our undergraduate curriculum: Our students are often familiar with many of the graphs and analyses in Table 1, and thus our course exposes them again to these topics. However, by covering these topics in tandem, students often gain a deeper understanding of inferential tools by seeing how they operate graphically. Finally, the taxonomy emphasizes that graphs and inferential tools can and should complement each other, which encourages the three statistical principles mentioned at the beginning of this section.

### v2

@@ Abstract
Data visualization is a core part of statistical practice and is ubiquitous in many fields. Although there are numerous books on data visualization, instructors in statistics and data science may be unsure how to teach data visualization, because it is such a broad discipline. To give guidance on teaching data visualization from a statistical perspective, we make two contributions. First, we conduct a survey of data visualization courses at top colleges and universities in the United States, in order to understand the landscape of data visualization courses. We find that most courses are not taught by statistics and data science departments and do not focus on statistical topics, especially those related to inference. Instead, most courses focus on visual storytelling, aesthetic design, dashboard design, and other topics specialized for other disciplines. Second, we outline three teaching principles for incorporating statistical inference in data visualization courses, and provide several examples that demonstrate how to follow these principles. The dataset from our survey allows others to explore the diversity of data visualization courses, and our teaching principles give guidance for encouraging statistical thinking when teaching data visualization.
Keywords: course design, course survey, graphics, inference, statistics education
@@ 1 Introduction: How is Data Visualization Taught, and What Defining Role Can Statistics Play?
Nearly every field involves data visualization, where information from data are visualized in order to gain insight on some phenomena. Data visualizations can be as simple as a bar plot of categorical counts to as complex as an interactive dashboard involving maps, time series, networks, and other high-dimensional graphs. Given its ubiquity, data visualization is taught across many fields, each with their own unique goals. As a result, there are many books dedicated to data visualization within specific fields, including business (Soukup and Davidson, , 2002; Knaflic, , 2015; Zheng, , 2017), journalism (Cairo, , 2012), and the humanities (Engebretsen and Kennedy, , 2020). There are also many books about data visualization in general (Tufte, , 2001; Few, , 2004; Chen et al., , 2007; Wilkinson, , 2012; Wilke, , 2019; Cairo, , 2016, 2019; Schwabish, , 2021; Cairo, , 2023; Yau, , 2024; Unwin, , 2024) and data visualization software (Sarkar, , 2008; Jones, , 2014; Wickham, , 2010, 2016; Murray, , 2017; Wickham et al., , 2019; Sievert, , 2020). Data visualization is such a pervasive and diverse area that, perhaps paradoxically, instructors in statistics, data science, and related fields may be unsure how to teach data visualization.
Indeed, the importance of data visualization in the practice and education of statistics has been emphasized for decades. For example, in the 2016 Guidelines for Assessment and Instruction in Statistics Education (GAISE) College Report endorsed by the American Statistical Association, one of the 10 goals is that students “should be able to produce graphical displays and numerical summaries and interpret what graphs do and do not reveal” (Carver et al., , 2016). More generally, data visualization has been important for knowledge generation and communication for centuries (for historical reviews of data visualization, see, e.g., Tufte, 2001 and Friendly, 2008). Maps for navigation were arguably the earliest form of data visualization; in the 1600s, visuals were used to understand variations in scientific measurements, which are often considered the first statistical visualizations; and in the 1800s, data visualizations were novelly used to argue for social and political change by, e.g., Florence Nightingale and W.E.B. Du Bois. Although data visualization is likely used in most (if not all) introductory statistics courses, instructors and departments may wonder if there is value in teaching data visualization specifically, perhaps even in a standalone course. For example, our statistics and data science department at Carnegie Mellon University offers data visualization courses at the undergraduate and graduate level. And even if departments do not plan to dedicate a semester-long class to data visualization, they may wonder how they can incorporate instruction of data visualization into existing courses.
@@ 2. Question 2: What defining role can statistics play in teaching data visualization?
To answer the first question, we collected data on 135 highly ranked colleges and universities, and determined which courses on data visualization, if any, were offered at each school in the 2022-2023 academic year. Among these 135 schools, we identified 270 data visualization courses. We recorded the name of each course, whether it was an undergraduate- and/or graduate-level course, the department(s) that offered it, and the topics and software taught based on its course description. There were three notable conclusions from our search. First, many schools offered data visualization courses (94 out of 135, or 69.6%), and often several departments within a school offered their own data visualization courses. The most frequent departments were related to computer science, data, information, and management, but we also found courses in less stereotypically data-centric fields, like geography, journalism, and the visual arts. Second, we found that most data visualization courses were not taught in statistics and data science departments (220 out of 270, or 81.5%), which may come as a surprise to some readers. In part this is because we deemed a course as a data visualization course only if at least half of the course’s material seemed dedicated to teaching data visualization itself. We found that while data visualization is frequently taught as one component of statistics courses (e.g., exploratory data analysis in introductory statistics courses), it is also rarely the focus of the course. Thus, introductory statistics courses were often not included in our survey. Finally, we found that most data visualization courses did not focus on statistical topics, and instead tended to specialize in visual storytelling, aesthetic design, or software use and development. In particular, the vast majority of courses did not focus on key aspects of statistical thinking: testing, uncertainty quantification, and modeling. For example, although many courses taught students how to visually examine trends and patterns in data, they rarely taught students how to visualize uncertainty in these trends or determine if these were “statistically significant” findings.
@@
To demonstrate how statistics departments can do this, we discuss how data visualization is taught in our statistics and data science department at Carnegie Mellon University, both in a 100-student semester-long undergraduate course (which serves many majors) as well as a 30-student half-semester master’s course. When our department first started teaching data visualization, it focused on visual design choices and how to use statistical software, which are common topics covered by many other college-level courses, as revealed by our survey. But over time, our course has evolved to put more emphasis on statistical issues that arise when creating data visualizations. For example, how should students pinpoint whether a graphical finding is “statistically significant,” and how can they quantify their uncertainty about trends and patterns they visualize? Furthermore, if multiple comparisons can be made within a single graph, how should the visual account for these multiple comparisons in order for inference to be correct? Addressing these questions forced us to develop a taxonomy of graphs and analyses, such that students can better understand how inferential tools can be properly used to complement their graphs, and vice versa. Our course isn’t the only one that emphasizes statistical aspects of data visualization; in our survey, we found several other courses that emphasize similar topics. Nonetheless, these courses are by far the minority among data visualization courses.
@@ 2.1 Description of Survey
For each university and college, we searched their online course catalog for credit-granting courses in the 2022-2023 academic year that contained one or more of the following words or combinations of words in the course title or description: (1) data visualization, (2) visualization / visualize / visualizing, (3) graph(s) / graphics / graphing, (4) plot(s) / plotting, and/or (5) chart(s) / charting. After identifying all courses that contained these words, we carefully reviewed the course titles, descriptions, and syllabi or websites (where available), and declared a course as a data visualization course only if its primary focus was indeed data visualization. By “primary focus,” we mean that at least half of the course’s material seemed dedicated to teaching data visualization itself. This criterion was meant to differentiate between courses whose primary motivation was teaching how to make effective data visualizations and courses that only used visualization in service of teaching a different concept. For instance, there are many introductory statistics and data science courses that cover data visualization in some capacity, e.g., using exploratory data analysis to describe a sample population, or using graphs to assess the plausibility of assumptions used for inference. However, in these classes, data visualization is usually only restricted to a small portion of the course, or visualization is only used to illustrate a different concept that was the main motivation for the course. Thus, we often did not consider introductory statistics courses as data visualization courses in our survey, unless their primary focus was indeed data visualization. For simplicity, we excluded courses in professional programs (medical school, nursing school, law school, and MBA programs), such that we focus on general undergraduate and graduate programs. Additionally, we only considered a course a data visualization course if its instruction on visualization was primarily data-centric, and not just visualization in general. For example, there are graphics courses that focus on creating video game graphics, which is a different context than visual representations of data that we are interested in. Similarly, there are many courses that focus on dashboard design or geospatial software, but only some of these involve data; we only considered these courses as data visualization courses if they indeed focused primarily on visualizing information from data (e.g., data dashboards or data-based maps).
@@
For each course that fit our search criteria, we gathered several pieces of information relevant to understanding the different contexts in which data visualization is taught. This includes the academic level (undergraduate, graduate, or both) and department in which the course was taught, as well as what software was used if listed in the course description or available syllabus/website. To additionally gauge the level of statistical content included in data visualization courses, we tracked if any of the following topics were mentioned in the course’s description, syllabus, or website (if available): hypothesis testing, confidence intervals, statistical modeling (e.g., linear regression), spatial data, clustering, time series, high dimensional analyses (e.g., principal component analysis), text data, networks, and interactive graphics (e.g., Shiny, D3.js). When possible, we also included a publicly available URL that was associated with each course, which others can use to inspect the content of a course. Course-specific URLs were available for 166 (61.5%) of courses.
@@
All replication materials for our survey results, including the dataset of data visualization courses we identified in our survey, are available at github.com/ryurko/teaching-data-viz. In what follows, we present several results about this dataset, thereby describing the current landscape of data visualization courses at U.S. colleges and universities. We hope that this dataset can also serve as a useful resource for instructors who want to glean further insights about data visualization courses, or use it as an example dataset in their own courses.
@@ 2.2 Survey Results
In total, we identified 270 data visualization courses across this pool of universities/colleges. Figure 1 displays the number of data visualization courses across schools; most schools have fewer than five courses, with a right-skew denoting places that offer several courses. Georgetown University (15 courses), New York University (14), and Northeastern University (13) stand out with the highest numbers.
@@
Figure 1: Number of data visualization classes taught within a school.
@@
One of our main interests is understanding how many data visualization courses are taught by statistics and data science departments. Among the 270 courses, there are 163 unique department names, making it difficult to categorize courses by department. Indeed, departments ranged from quantitative fields (e.g., computer science, economics, engineering, mathematics, statistics) to social sciences (e.g., education, psychology, sociology) to professional domains (e.g., analytics, business, government, journalism) to humanities and the arts (e.g., anthropology, design, history, visual arts, writing), demonstrating the ubiquity of data visualization. Figure 2 shows a word cloud of the most common words across department names among these 270 courses; when creating this word cloud, we removed stop words (such that common words like “of” and “the” were removed) and performed stemming (such that words like “science” and “sciences” were combined as “scienc”). Only words that were used at least three times across department names are displayed in Figure 2. Broadly speaking, the most common departments are related to computer science, data, information, and management. We also see many departments related to analytics, business, engineering, public policy, and statistics. Finally, the diversity of less common departments is notable, including the arts, biology, geography, health, journalism, and mathematics.
@@
To understand the involvement of statistics and/or data science departments specifically, we identified department names that included either “stat” or “data” in their name. [^2] The only departments that mentioned “data” but not “data science” mentioned “data analytics;” the only exception was a department called “Data for Political Research.” In order to take an inclusive view of data science, we also considered these as data science departments. We found 50 of the 270 courses (18.5%) were taught by statistics and/or data science departments. This proportion was also fairly consistent across the level of courses: Figure 3 displays the number of data visualization courses at the undergraduate or graduate level, or both (e.g., cross-listed courses), colored by whether the course’s department was in statistics or data science. A roughly equal number of data visualization courses are offered at the undergraduate or graduate level, where 21.7% of undergraduate courses, 18.8% of graduate courses, and 12.7% of cross-listed courses are taught by statistics or data science departments. Taken together, Figures 2 and 3 demonstrate that many fields teach data visualization, most of which are outside of statistics and data science.
@@
Figure 2: Word cloud displaying the top words in the department names for the 270 identified data visualization courses. All words displayed were used at least three times across department names.
@@
Meanwhile, Figure 4(b) shows the number of topics that are taught across courses, and we see that the vast majority of courses (226, or 88.3%) teach two or fewer of these 10 topics. For example, we found that many courses that taught interactive graphics focused exclusively on how to create data visualization dashboards, and many courses that taught spatial data focused exclusively on how to make data-based maps with geospatial software. The only course that covers nine topics is our own course in the Department of Statistics and Data Science at Carnegie Mellon University (which we discuss in the next section), and the only course that covers eight topics is from the Department of Statistics at Colby College. [^3] The only topic our course doesn’t cover is networks. Notably, the instructor who developed the Colby College course was previously a PhD student in our department and taught our data visualization course. However, we want to emphasize that this does not suggest that courses are inadequate by not covering these topics; many courses are certainly teaching other useful content, e.g., visual storytelling and design. Rather, departments understandably focus on topics most related to their own fields, instead of teaching statistical topics, especially those related to inference.
@@
We acknowledge that there are limitations with this survey. First, there are likely other data visualization courses not captured by our survey (e.g., at other schools not considered here, courses that did not mention keywords we searched for, or courses whose descriptions or syllabi are not reflective of the data visualization content taught in the course). Similarly, some courses may cover statistical topics but did not mention them on publicly available sources we could access. Finally, this survey does not capture all possible topics taught in data visualization courses, because we focused on statistical topics. Nonetheless, this survey demonstrates that data visualization is taught across many colleges and universities, and by a broad range of disciplines that are largely outside of statistics or data science. Furthermore, most data visualization courses do not emphasize key aspects of statistical thinking (testing, uncertainty quantification, and modeling), and instead focus on topics unique to particular disciplines (e.g., storytelling, visual design, dashboard design, or specialized software). Thus, to complement current offerings in data visualization, in the following section we demonstrate how instructors can incorporate statistical thinking when teaching data visualization.
@@ 3.1 Description of Our Course, and a Taxonomy of Graphs
Our undergraduate data visualization course in the Department of Statistics and Data Science at Carnegie Mellon University (CMU), titled Statistical Graphics and Visualization, has existed for almost 20 years and has evolved over time. Currently, the course is offered every fall and spring semester, and serves approximately 100 students each semester. The prerequisite is two semesters of introductory statistics (not calculus-based), such that students have a basic understanding of hypothesis testing, statistical models (including linear regression), and how to apply them to real datasets. We use R to implement graphs and analyses, but students are not required to have prior programming experience. In addition to counting as an elective for our statistics and data science majors, the course also counts towards our college’s general education requirement, specifically the “design” requirement, which consists of courses focused on designing products to solve problems. Our department is within CMU’s Dietrich College of Humanities and Social Sciences, and we regularly teach majors in economics, English, history, information systems, psychology, social and decision sciences, and other fields, in addition to our own statistics and data science majors. Thus, our course is meant for a broad group of students from many disciplines who are at a sophomore-level or above in terms of their statistical background.
@@
Histograms, smoothed densities,
empirical CDFs
@@
Some readers may wonder how the taxonomy in Table 1 is related to the grammar of graphics framework, which is a common conceptual model for organizing, implementing, and teaching data visualization (Wickham, , 2010; Wilkinson, , 2012). In short, the grammar of graphics framework outlines seven components that map data to aspects of a graph; this includes how data are mapped to the geometry, coordinate system, and aesthetics of a graph. Table 1 includes two of these components: the variables and statistics that are used for a graphic. We view the grammar of graphics framework and the taxonomy in Table 1 as complementary to each other: The grammar of graphics framework formulates how graphs themselves are constructed, and Table 1 formulates how prototypical graphs are related to statistical inference. This way, students can more easily understand how graphs fit into a statistical inference paradigm.
Because of this, we’ve found that there are several pedagogical benefits to organizing the first half of the course with the taxonomy in Table 1. First, students are typically quite good at determining whether a variable is categorical or quantitative, but they may find it challenging to determine which graphs or analyses are most appropriate for a given dataset. Thus, organizing by variable type gives students a straightforward scaffolding for navigating graphs and analyses. Second, it prepares students to make and interpret more complex graphs in the second half of the course (e.g., high-dimensional visualizations, maps, interactive graphs), which frequently use the graphs in Table 1 as building blocks. Third, the taxonomy encourages a spiral learning approach in our undergraduate curriculum: Our students are often familiar with many of the graphs and analyses in Table 1, and thus our course exposes them again to these topics. However, by covering these topics in tandem, students often gain a deeper understanding of inferential tools by seeing how they operate graphically. Finally, the taxonomy emphasizes that graphs and inferential tools can and should complement each other, which encourages the three statistical principles mentioned at the beginning of this section.